Exercise 16-30 (Algo) Sales Activity Variance (LO 16-3) The following data are available for the most recent year of operations for Slacker & Sons. The revenue portion of the sales activity variance is $288,000 F. Master budget based on actual

The following data are available for the most recent year of operations for Slacker & Sons. The revenue portion of the sales activity variance is $288,000 F.
 





Master budget based on actual sales of 161,000 units:


Revenue$3,600,000
Materials
861,000
Labor
636,000
Variable manufacturing overhead and administrative costs
136,000
Fixed manufacturing overhead and administrative costs
410,000

 

Required:

a. How many units were actually sold in the most recent period?

b. Prepare a sales activity variance for the most recent year for Slacker & Sons.
Answer #1

Master budget based on actual sales of 161,000 units:   

Revenue                                                                                $3,600,000 

Materials                                                                                  861,000 

Labor                                                                                          636,000 

Variable manufacturing overhead and administrative costs       136,000 

Fixed manufacturing overhead and administrative costs             410,000 


A) 288,000/ 3,600,000= 0.08 +1= 1.08     just add by 1 

1.08* 161,000= 173,800 Actual sales (answer)


B) 

Master budget column:

sales revenue    3,600,000

materials       861,000

direct labor      636,000

variable manufacturing and administrative      136,000

total variable cost   861,000+ 636,000+ 136,000=  1,633,000

contribution margin     3,600,000- 1,633,000= 1,967,000

Manufacturing overhead and administrative   410,000

total fixed cost    410,000

operating profits  1,967,000- 410,000= 1,557,000


sales activity variance column:

sales revenue  280,000 F

materials   861,000* 8% (or 0.08 from A)= 68,880 U

direct labor  8% *636,000= 50,880 U

variable manufacturing and administrative  8%* 136,000= 10,880 U 

total variable cost   8% * 1,633,000= 130,640 U

contribution margin 8%* 1,967,000= 157,360 F

Manufacturing overhead and administrative 0

total fixed cost    0

operating profits 157,360 F

                      

flexible budget column:

sales revenue 3,660,000+ 288,000= 3,888,000

materials   861,000+ 68,880= 929,880

direct labor 636,000+ 50,880= 686,880

variable manufacturing and administrative  136,000+ 10,880= 146,880   

total variable cost 929,880+ 686,880+ 146,880= 1,763,640

contribution margin  3,888,000+ 1,763,640= 2,124,360

Manufacturing overhead and administrative 410,000

total fixed cost  410,000

operating profits 2,124,360- 410,000= 1,714,360

select